A jet travels 4912 miles against the wind in 8 hours and 6192 miles with the wind in the same amount of time. What is the rate of the jet in still air and what is the rate of the wind?
Answer by mananth(16946)

Plane speed =x mph
wind speed =y mph
against wind 8 hours
with wind 8 hours

Distance with wind 6192 miles distance against wind 4912
t=d/r against wind (x-y)
4912.00 / ( x - y )= 8.00

8.00 x - -8.00 y = 4912.00 ....................1
with wind (x+y)
6192.00 / ( x + y )= 8.00
8.00 ( x + y ) = 6192.00
8.00 x + 8.00 y = 6192.00 ...............2
Multiply (1) by 1.00
Multiply (2) by 1.00
we get 2.00
8.00 x + -8.00 y = 4912.00
8.00 x + 8.00 y = 6192.00
16.00 x = 11104.00
/ 16.00
x = 694 mph

plug value of x in (1) y
8 x -8 y = 4912
5552 -8 -5552 = 4912
-8 y = 4912
-8 y = -640 mph
y = 80
Plane 694 mph
wind 80 mph